# The discrete nonlinear Schrödinger equation with linear gain and nonlinear loss: the infinite lattice with nonzero boundary conditions and its finite dimensional approximations

### v1

@@ III The case of the lattice with periodic boundary conditions
    $\boldsymbol{\omega}(w_{n}(0))=\gamma_{*},$ (3.6)
    where $\gamma_{*}$ is the time-periodic orbit defined by the only non-trivial plane wave solution of constant amplitude $A_{*}$ with frequency $\tilde{\omega}$ and wave number $q$ satisfying (3.4), when we set in (3.4) $A(t)\equiv A_{*}$, that is,
@@ 2. We start by seeking solutions of the form
Now, we may use (3.21) in the solution (3.3) and (3.27)-(3.28) (or alternatively (3.18) and (3.24)), to see that as $t\rightarrow\infty$, $\gamma_{*}$ is the orbit defined by the solution
@@
The main consequence of Theorem II.1 is that the dynamical system defined by the periodic lattice (1.1)-(1.3) is globally asymptotically stable for any initial condition in $\ell^{2}_{per}$: for arbitrary initial conditions all the orbits are converging to a state whose averaged norm satisfies the bound (3.2). In the special case where the initial condition is defined by a solution of the form (3.3), the attractor is trivial, defined by the only plane wave solution which can have constant amplitude, which is $A^{*}$. This global asymptotic stability do not excludes the emergence of instability effects. Nevertheless, these effects can be only transient, prior the convergence of the system to its ultimate state. We investigate the potential of these transient instability effects in the framework of the linear modulation instability analysis of plane waves. The motivation is that the fate of such initial conditions is predicted by the Theorem II.1 as stated above.
The modulational instability (MI) of plane-waves is well studied in the case of the Hamiltonian DNLS [19]. We will perform here this analysis for the system (1.1) with the presence of the linear gain and nonlinear loss effects. The significance of these findings lies on the fact that they prove that the simplest initial data of the form (3.31) have interesting dynamics prior the convergence to the attractor, and that even this attractor can be transiently, modulationally unstable.
@@
The linear homogeneous system (3.34) for $(b_{0},\psi_{0})$ has a nontivial solution if and only if its determinant is zero. This request provides the modulational instability formula:
$\Lambda^{2}-2\mathrm{i}\delta A^{2}\Lambda-\Gamma\left(\Gamma-2A^{2}\right)=0,$ (3.35)
with $\Lambda=\Omega-2k\sin(hQ)\sin(hq)$ and $\Gamma=4k\sin^{2}(Q/2)\cos(hq)$. We elaborate further on the formula (3.35), to derive the conditions for the modulation instability of the examined plane wave solution. Since (3.35) is a quadratic equation in $\Lambda$ with complex coefficients, it has only complex roots. The discriminant of (3.35) is
@@
We will illustrate the above transient modulation instability of the plane wave attractor with numerical simulations, in the next section.
@@
This section is devoted to the first part of the numerical investigations of the paper. We illustrate the transient modulation instability of the plane wave attractor, which is verified by the modulation instability analysis of the section III.1. For this purpose, we study the dynamics of the simplest plane wave initial conditions, defining perturbations of the plane wave attractor,
@@
Due to the periodic boundary conditions (1.3) and discreteness, the physically meaningful $K\in[0,N/2]$, $K\in\mathbb{N}$. In terms of the modulation instability analysis of section III.1, $q=\pi K/L$, and the modulation instability condition (3.38) is satisfied for $250<K<$ 25, defining the unstable modes. The rest of the modes $K$ are modulationally stable.
@@
Figure 1: Convergence to the plane wave attractor with amplitude $A_{*}=1$ for the stable mode $K=45$. Rest of parameters: $L=50$, $h=1$, $\gamma=-\delta=1.5$. Panel (a): Fourier spectrum of the initial condition (4.1) for $A_{p}=2$ and $A_{p}=-0.999$. Panel (b): Fourier spectrum of the solutions at $t=5$, for $A_{p}=2$ and $A_{p}=-0.999$. Panel (c): Convergence of the orbits of the central node of the lattice $u_{0}(t)$ in the phase plane $\big{(}\textrm{Re}(u_{0}(t)),\,\textrm{Im}(u_{0}(t))\big{)}$ to the limit cycle of radius $A_{*}=1$; internal (red) orbit for $A_{p}=-0.999$ and external (blue) orbit for $A_{p}=2$. Details are given in the text.
The first example concerns the dynamics of the stable modes. We choose the stable mode $K=45$. According to the results of Section III, for arbitrary perturbations $A_{p}$, the orbit starting from the initial condition (4.1) will converge to the plane wave attractor with amplitude $A_{*}=1$ and the initially chosen mode $K=45$. The dynamics for the amplitude perturbations $A_{p}=2$ and $A_{p}=-0.999$ is depicted in Fig. 1. Panel (a) depicts the Fourier spectrum of the initial condition (4.1) for $A_{p}=2$ and $A_{p}=-0.999$, where $K=45$ is the chosen stable mode. Panel (b) depicts the Fourier spectrum of the solutions at $t=5$ where convergence to the attractor is attained, for both cases of the perturbations $A_{p}$. It shows that the attractor preserves the initial wave number $K=45$, as predicted. No other modes are excited transiently (intermediate snapshots are omitted as they are identical to (b)), since the initial mode is stable. Panel (c) illustrates the convergence to the attractor in the plane $\big{(}\mathrm{Re}(u_{0}(t)),\mathrm{Im}(u_{0}(t))\big{)}$ which is the phase plane for the central node $u_{0}(t)$ of the lattice located at $n=0$. The dynamics is portrayed as a convergence to the asymptotically stable limit cycle of radius $A_{*}$; the orbit in the interior of the limit cycle is the one for $A_{p}=-0.999$ and the orbit in the exterior of the limit cycle is the one for $A_{p}=2$.
@@
The second example concerns the case of the unstable modes. We choose the unstable mode $K=8$. According to the results of Section III.1 the corresponding orbit should converge to the plane wave attractor with amplitude $A_{*}=1$. Concerning the wave number of the attracting state, the analytical arguments provide the prediction for its choice: For the linear gain strength $\gamma=1.5$ the convergence to the amplitude $A_{*}$ should be fast. On the other hand, the initial unstable wave number should introduce transient modulation instability effects. While the convergence to the steady-state amplitude $A_{*}$ has a fast exponential rate, the emergence of instability effects will occur later, perturbing the phase of the attracting state, transiently. Nevertheless, the ultimate state should choose a stable wave number, since it is the globally asymptotically stable state for the system. This prediction is fully illustrated in Figures 2 and 3 depicting the dynamics for $A_{p}=2$. Figure 2 shows the evolution of the Fourier spectrum of the solution. The initial wave number $K=8$, of panel (a) is preserved as shown in panel (b) for $t=5$; at this time the system has almost converged to the plane wave with amplitude $A_{*}=1$ at a fast exponential rate as shown in the panel (a) of Figure 3, and MI effects are not yet visible. These are developed later as depicted in panel (c) of Figure 2, demonstrating the excitation of the whole Fourier spectrum at $t=125$. Panel (b) of Figure 3 portrays the deformed orbit for $t\in[0,125]$ which is a time period within the stage of modulation instability. The instability can last for large times, however, eventually, the global attractor is attained and its wave number should be selected by the stability band. This is exactly the case, depicted in panel (d) of Figure 2, showing the snapshot of the Fourier spectrum at $t\in[3600,3700]$. At this time the only active wavenumber is the stable mode $K=44$ which is the one choosed by the plane wave attractor of amplitude $A_{*}$.
@@
V Numerical Study II: Simulating the infinite lattice by its finite lattice approximations
@@
and the fact that any plane wave initial condition is attracted by the orbit $\gamma_{*}$.
@@
The transient dynamics prior to the convergence to $\gamma_{*}$ is illustrated. This involves the excitation of the full spectrum at the early stages of the evolution as it is shown in panel (b) for $t=40$, while progressively, the active Fourier modes are reduced, as it is shown in the snapshots (c) for $t=80$, (d) for $t=200$ and (e) for $t=600$. When the orbit $\gamma_{*}$ is reached, the only active mode is the stable mode $K=50$, see panel (f). The same dynamics was observed for other choices of the lattice parameters and gain and loss strengths.
@@
For the periodic lattice (1.1)-(1.3) the numerical simulations illustrate that the dynamics prior the convergence to the global attractor are associated with the emergence of characteristic patterns of MI dynamics and the emergence of extreme wave events. This is a consequence of the transient modulation instability discussed above. Our numerical investigations consider two examples of initial conditions of the form (5.2), with different rates of localization. The first concerns (5.2) with the quadratically decaying function $f(x)$ (5.3), and the second with the exponentially decaying function
@@
Figure 5: Top row: Spatiotemporal evolution of the density $|u_{n}(t)|^{2}$ for the DNLS equation (1.1), with initial conditions (5.2) for $A=A_{*}=0.5$ when $\gamma=0.0025$ and $\delta=-0.01$. Panel (a) for $f(x)$ (5.3) with $\lambda_{1}=\lambda_{2}=1$ and $\lambda_{3}=4$. Panel (b) for $f(x)$ (5.4) with $\sigma=0.6$, $\rho=1$. Bottom row: The spatiotemporal evolution of the density $|\phi_{n}(t)|^{2}$ of the solutions of the AL-lattice (5.6). Panel (c) for $f(x)$ (5.3) and panel (d) for $f(x)$ (5.4) with parameters as above. Rest of the lattice parameters for both systems: $L=200$, $h=1$. More details in the text.
@@
Figure 7: Spatiotemporal evolution of the density $|u_{n}(t)|^{2}$ for the DNLS equation (1.1) with initial conditions (5.2) with $A=0.5$ when $\gamma=0.01$, $\delta=-0.01$, i.e., $A\neq A_{*}$ Panel (a) for $f(x)$ (5.3) with $\lambda_{1}=\lambda_{2}=1$ and $\lambda_{3}=4$. Panel (b) for $f(x)$ (5.4) with $\sigma=0.6$, $\rho=1$. Rest of the lattice parameters for both systems: $L=200$, $h=1$. More details in the text.
@@
the rational analytical solution of the integrable Ablowitz–Ladik (AL) lattice
$\mathrm{i}\dot{\phi}_{n}+k(\phi_{n+1}-\phi_{n}+\phi_{n-1})+|\phi_{n}|^{2}(\phi_{n-1}+\phi_{n+1})=0,$ (5.6)
@@
The patterns of the DNLS (1.1) share major characteristics of the ones emerging in the dynamics the AL- lattice when the same initial conditions are used, as it is shown in the bottom row of Figure 5, where panel (c) corresponds to the initial condition with $f(x)$ (5.3) and panel (d) to the initial condition with $f(x)$ (5.4). In particular, the patterns of the AL system are specific examples, here in the case of the integrable AL-lattice, of the universal behavior of modulationally unstable media [13]. Let us recall that in [13], the long-standing open question about the nonlinear stage of modulation instability on the infinite line for the integrable NLS equation
@@
was fully resolved. It was proved in [14], that for any given initial condition $u(x,0)$ representing a sufficiently localized initial perturbation of the constant background $A$, the solution $u(x,t)$ of (5.7) is $u(x,t)=u_{asymp}(x,t)+\mathrm{O}(1/\sqrt{t})$ as $t\rightarrow\infty$. Cruciallly, $u_{asymp}(x,t)$ has different forms in different sectors of the $xt$-plane, namely: (i) in the two “plane-wave” regions, $x<-4\sqrt{2}At$ and $x>4\sqrt{2}At$, we have that $|u_{asymp}(x,t)|=A$, that is, the solution has the same amplitude as the undisturbed background $A$ (ii) in the “modulated elliptic wave” region $-4\sqrt{2}At<x<4\sqrt{2}At$ the solution is expressed by a slow modulation of the elliptic solutions of (5.7). In [13], analytical arguments corroborated with numerical simulations provided evidence that this behavior is universal, as it is shared by a wide class of NLS-type equations and systems, including the AL-lattice (5.6).
@@
In this context, of the universality of the MI behavior, we claim that the findings depicted for the DNLS (1.1) in Figure 5 are important. First we should stress that the DNLS (1.1) is a system at least “two steps forward” regarding the breaking of the integrability barrier defined by the AL system, since it is a dissipative perturbation of the non-integrable Hamiltonian DNLS. Second, the theoretical results of Theorems II.1 and III.1 accompanied with the MI analysis of the plane-wave of amplitude $A_{*}$ are fully explaining the dynamics observed in panels (a) and (b). The illustrated dynamics show that the breaking of integrability induced by the DNLS (1.1) is not dramatic. It shares the major characteristics of the dynamics of the integrable AL-lattice, that is, the structure comprised of the two outer, quiescent sectors separated by the wedge-shaped central region characterized by the oscillatory behavior. In the outer sectors, the amplitude of $u_{n}(t)$ is $A_{*}$, of the undisturbed background, the only one which may sustain localized solutions in the case of the infinite lattice. On the other hand, since the numerical solutions are generated by solving numerically the periodic lattice, this background will never lose stability in terms of the amplitude in the outer regions since the solutions are initiated on the stable attractor of the finite system regarding its amplitude. Yet for the finite dimensional system, the MI dynamics in the wedge-shaped central region can be only transient as it is proved in Theorem III.1 and analyzed in the discussion of section V.1. Importantly, in the case of the finite lattice the wedge will eventually disappear as the system will select a stable wave number defining the global attractor, and the asymptotic state of the solution will be the plane wave of amplitude $A_{*}$ with the stable wavenumber.

### v3

@@ III The case of the lattice with periodic boundary conditions
    $\boldsymbol{\omega}(w_{n}(0))=\mathcal{C}_{*},$ (3.6)
    where $\mathcal{C}_{*}$ is the time-periodic orbit defined by the only non-trivial plane wave solution of constant amplitude $A_{*}$ with frequency $\tilde{\omega}$ and wave number $q$ satisfying (3.4), when we set in (3.4) $A(t)\equiv A_{*}$, that is,
@@ 2. We start by seeking solutions of the form
Now, we may use (3.21) in the solution (3.3) and (3.27)-(3.28) (or alternatively (3.18) and (3.24)), to see that as $t\rightarrow\infty$, $\mathcal{C}_{*}$ is the orbit defined by the solution
@@
The main consequence of Theorem II.1 is that the dynamical system defined by the periodic lattice (1.1)-(1.3) is globally asymptotically stable for any initial condition in $\ell^{2}_{per}$: for arbitrary initial conditions all the orbits are converging to a state whose averaged norm satisfies the bound (3.2). In the special case where the initial condition is defined by a solution of the form (3.3), the attractor is trivial, defined by the only plane wave solution which can have constant amplitude, which is $A^{*}$. This global asymptotic stability do not excludes the emergence of instability effects. Nevertheless, these effects can be only transient, prior the convergence of the system to its ultimate state. We investigate the potential of these transient instability effects in the framework of the linear MI analysis of plane waves. The motivation is that the fate of such initial conditions is predicted by the Theorem II.1 as stated above.
The modulation instability of plane-waves is well studied in the case of the Hamiltonian DNLS [26]. We will perform here this analysis for the system (1.1) with the presence of the linear gain and nonlinear loss effects. The significance of these findings lies on the fact that they prove that the simplest initial data of the form (3.31) have interesting dynamics prior the convergence to the attractor, and that even this attractor can be transiently, modulationally unstable.
@@
The linear homogeneous system (3.34) for $(b_{0},\psi_{0})$ has a nontivial solution if and only if its determinant is zero. This request provides the modulation instability formula:
$\Lambda^{2}-2\mathrm{i}\delta A_{*}^{2}\Lambda-\Gamma\left(\Gamma-2A_{*}^{2}\right)=0,$ (3.35)
with $\Lambda=\Omega-2k\sin(hQ)\sin(hq)$ and $\Gamma=4k\sin^{2}(Q/2)\cos(hq)$. We elaborate further on the formula (3.35), to derive the conditions for the MI of the examined plane wave solution. Since (3.35) is a quadratic equation in $\Lambda$ with complex coefficients, it has only complex roots. The discriminant of (3.35) is
@@
We will illustrate the above transient MI of the plane wave attractor with numerical simulations, in the next section.
@@
This section is devoted to the first part of the numerical investigations of the paper. We illustrate the transient MI of the plane wave attractor, which is verified by the analysis of the section III.1. For this purpose, we study the dynamics of the simplest plane wave initial conditions, defining perturbations of the plane wave attractor,
@@
Due to the periodic boundary conditions (1.3) and discreteness, the physically meaningful $K$ are integers in the interval $[0,N/2]$. In terms of the analysis of section III.1, $q=\pi K/L$, and the MI condition (3.38) is satisfied for $250<K<$ 25, defining the unstable modes. The rest of the modes $K$ are modulationally stable.
@@
Figure 1: Convergence to the plane wave attractor with amplitude $A_{*}=1$ for the stable mode $K=45$. Rest of parameters: $L=50$, $h=1$, $\gamma=-\delta=1.5$. Panel (a): Fourier spectrum of the initial condition (4.1) for $A_{p}=2$ and $A_{p}=-0.999$. Panel (b): Fourier spectrum of the solutions at $t=10$, for $A_{p}=2$ and $A_{p}=-0.999$. Panel (c): Convergence of the orbits of the central node of the lattice $u_{0}(t)$ in the phase plane $\big{(}\textrm{Re}(u_{0}(t)),\,\textrm{Im}(u_{0}(t))\big{)}$ to the limit cycle of radius $A_{*}=1$; internal (red) orbit for $A_{p}=-0.999$ and external (blue) orbit for $A_{p}=2$. Details are given in the text.
The first example concerns the dynamics of the stable modes. We choose the stable mode $K=45$. According to the results of Section III, for arbitrary perturbations $A_{p}$, the orbit starting from the initial condition (4.1) will converge to the plane wave attractor with amplitude $A_{*}=1$ and the initially chosen mode $K=45$. The dynamics for the amplitude perturbations $A_{p}=2$ and $A_{p}=-0.999$ is depicted in Fig. 1. Panel (a) depicts the Fourier spectrum of the initial condition (4.1) for $A_{p}=2$ and $A_{p}=-0.999$, where $K=45$ is the chosen stable mode. Panel (b) depicts the Fourier spectrum of the solutions at $t=10$ where convergence to the attractor is attained, for both cases of the perturbations $A_{p}$. It shows that the attractor preserves the initial wave number $K=45$, as predicted. No other modes are excited transiently (intermediate snapshots are omitted as they are identical to (b)), since the initial mode is stable. Panel (c) illustrates the convergence to the attractor in the plane $\big{(}\mathrm{Re}(u_{0}(t)),\mathrm{Im}(u_{0}(t))\big{)}$ which is the phase plane for the central node $u_{0}(t)$ of the lattice located at $n=0$. The dynamics is portrayed as a convergence to the asymptotically stable limit cycle of radius $A_{*}$; the orbit in the interior of the limit cycle is the one for $A_{p}=-0.999$ and the orbit in the exterior of the limit cycle is the one for $A_{p}=2$.
@@
The second example concerns the case of the unstable modes. We choose the unstable mode $K=8$. According to the results of Section III.1 the corresponding orbit should converge to the plane wave attractor with amplitude $A_{*}=1$. Concerning the wave number of the attracting state, the analytical arguments provide the prediction for its choice: For the linear gain strength $\gamma=1.5$ the convergence to the amplitude $A_{*}$ should be fast. On the other hand, the initial unstable wave number should introduce transient MI effects. While the convergence to the steady-state amplitude $A_{*}$ has a fast exponential rate, the emergence of instability effects will occur later, perturbing the phase of the attracting state, transiently. Nevertheless, the ultimate state should choose a stable wave number, since it is the globally asymptotically stable state for the system. This prediction is fully illustrated in Figures 2 and 3 depicting the dynamics for $A_{p}=2$. The dynamics for the corresponding unstable case of $A_{p}=-0.999$ is very similar and, thus, is not shown herein. Figure 2 shows the evolution of the Fourier spectrum of the solution. The initial wave number $K=8$, of panel (a) is preserved as shown in panel (b) for $t=5$; at this time the system has almost converged to the plane wave with amplitude $A_{*}=1$ at a fast exponential rate as shown in the panel (a) of Figure 3, and MI effects are not yet visible. These are developed later as depicted in panel (c) of Figure 2, demonstrating the excitation of the whole Fourier spectrum at $t=125$. Panel (b) of Figure 3 portrays the deformed orbit for $t\in[0,125]$ which is a time period within the stage of MI. The instability can last for large times, however, eventually, the global attractor is attained and its wave number should be selected by the stability band. This is exactly the case, depicted in panel (d) of Figure 2, showing the snapshot of the Fourier spectrum at $t\in[3600,3700]$. At this time the only active wavenumber is the stable mode $K=44$ which is the one selected by the plane wave attractor of amplitude $A_{*}$.
@@
V Numerical Study II: Simulating the infinite lattice with nonzero boundary conditions by the finite lattice with periodic boundary conditions
@@
and the fact that any plane wave initial condition is attracted by the orbit $\mathcal{C}_{*}$.
@@
The transient dynamics prior to the convergence to $\mathcal{C}_{*}$ is illustrated. This involves the excitation of the full spectrum at the early stages of the evolution as it is shown in panel (b) for $t=40$, while progressively, the active Fourier modes are reduced, as it is shown in the snapshots (c) for $t=80$, (d) for $t=200$ and (e) for $t=600$. When the orbit $\mathcal{C}_{*}$ is reached, the only active mode is the stable mode $K=50$, see panel (f). Qualitatively, the same dynamics was observed for other choices of the lattice parameters and gain and loss strengths.
We can summarize the results regarding the convergence of all initial data to the plane wave attractor as follows: Theorem III.1 and the MI analysis of section III.1 theoretically predict that the amplitude and spectrum of the initial condition will undergo a transient evolution before converging to the attractor. Ultimately, this process selects the amplitude $A_{*}$ and a wavenumber within the stable modes of the spectrum. It is important to note that this prediction is qualitative rather than quantitative, as the theoretical results do not specify the exact stable mode of the attractor.
@@
For the periodic lattice (1.1)-(1.3) the numerical simulations illustrate that the dynamics prior the convergence to the global attractor are associated with the emergence of characteristic patterns of MI dynamics and the emergence of extreme wave events. This is a consequence of the transient MI discussed above. Our numerical investigations consider two examples of initial conditions of the form (5.2), with different rates of localization. The first concerns (5.2) with the quadratically decaying function $f(x)$ (5.3), and the second with the exponentially decaying function
@@
Figure 5: Top row: Spatiotemporal evolution of the density $|u_{n}(t)|^{2}$ for the DNLS equation (1.1), with initial conditions (5.2) for $A=A_{*}=0.5$ when $\gamma=0.0025$ and $\delta=-0.01$. Panel (a) for $f(x)$ (5.3) with $\lambda_{1}=\lambda_{2}=1$ and $\lambda_{3}=4$. Panel (b) for $f(x)$ (5.4) with $\sigma=0.6$, $\rho=1$. Bottom row: The spatiotemporal evolution of the density $|\phi_{n}(t)|^{2}$ of the solutions of the AL-lattice (5.6). Panel (c) for $f(x)$ (5.3) and panel (d) for $f(x)$ (5.4) with parameters as above. Rest of the lattice parameters for both systems: $L=200$, $h=1$. The straight (black) lines depict the graphs of the lines $x=\pm 4\sqrt{2}At$. More details in the text.
@@
Figure 7: Spatiotemporal evolution of the density $|u_{n}(t)|^{2}$ for the DNLS equation (1.1) with initial conditions (5.2) with $A=0.5$ when $\gamma=0.01$, $\delta=-0.01$, i.e., $A\neq A_{*}$. Panel (a) for $f(x)$ (5.3) with $\lambda_{1}=\lambda_{2}=1$ and $\lambda_{3}=4$. Panel (b) for $f(x)$ (5.4) with $\sigma=0.6$, $\rho=1$. Rest of the lattice parameters for both systems: $L=200$, $h=1$. The straight (black) lines depict the graphs of the lines $x=\pm 4\sqrt{2}At$. More details in the text.
@@
the rational analytical solution of the integrable AL lattice
$\mathrm{i}\dot{\phi}_{n}+k(\phi_{n+1}-2\phi_{n}+\phi_{n-1})+|\phi_{n}|^{2}(\phi_{n-1}+\phi_{n+1})=0,$ (5.6)
@@
The patterns of the DNLS (1.1) share major characteristics of the ones emerging in the dynamics the AL- lattice when the same initial conditions are used, as it is shown in the bottom row of Figure 5, where panel (c) corresponds to the initial condition with $f(x)$ (5.3) and panel (d) to the initial condition with $f(x)$ (5.4). In particular, the patterns of the AL system are specific examples, here in the case of the integrable AL-lattice, of the universal behavior of modulationally unstable media [13]. Let us recall that in [13], the long-standing open question about the nonlinear stage of MI on the infinite line for the integrable NLS equation
@@
was fully resolved. It was proved in [21], that for any given initial condition $u(x,0)$ representing a sufficiently localized initial perturbation of the constant background $A$, the solution $u(x,t)$ of (5.7) is $u(x,t)=u_{asymp}(x,t)+\mathrm{O}(1/\sqrt{t})$ as $t\rightarrow\infty$. Cruciallly, $u_{asymp}(x,t)$ has different forms in different sectors of the $xt$-plane, namely: (i) in the two “plane-wave” regions, $x<-4\sqrt{2}At$ and $x>4\sqrt{2}At$, we have that $|u_{asymp}(x,t)|=A$, that is, the solution has the same amplitude as the undisturbed background $A$ (ii) in the “modulated elliptic wave” region $-4\sqrt{2}At<x<4\sqrt{2}At$ the solution is expressed by a slow modulation of the elliptic solutions of (5.7). In [13], analytical arguments corroborated with numerical simulations provided evidence that this behavior is universal, as it is shared by a wide class of NLS-type equations and systems, including the AL-lattice (5.6) and the Hamiltonian DNLS, which corresponds to the case of $\gamma=\delta=0$ of the dissipative DNLS (1.1) (see [13, Appendix C, pg. 903]). The findings of [13] regarding the universal behavior of MI for the DNLS systems, combined with the theoretical results proven herein, will serve as a roadmap for explaining our numerical results, as elaborated below.
@@
In this context, of the universality of the MI behavior, we claim that the findings depicted for the DNLS (1.1) in Figure 5 are important. First we should stress that the DNLS (1.1) is a system at least “two steps forward” regarding the breaking of the integrability barrier defined by the AL system, since it is a dissipative perturbation of the non-integrable Hamiltonian DNLS. Second, the theoretical results of Theorems II.1 and III.1 accompanied with the MI analysis of the plane-wave of amplitude $A_{*}$ are fully explaining the dynamics observed in panels (a) and (b) if combined with the findings of [13] mentioned above. The illustrated dynamics show that the breaking of integrability induced by the DNLS (1.1) is not dramatic. It shares the major characteristics of the dynamics of the AL-lattice and of the Hamiltonian DNLS observed in [13], that is, the structure comprised of the two outer, quiescent sectors separated by the wedge-shaped central region characterized by the oscillatory behavior. The straight (black) lines depict the graphs of the lines $x=\pm 4\sqrt{2}At$, which are the exact boundaries of the wedge in the case of the integrable NLS partial differential equation, calculated in [21]. It is crucial to remark that, according to the analysis of [13], the slopes of the boundaries are modified depending on the particular system considered; this modification also occurs in the case of the dissipative DNLS lattice (1.1). It is interesting to observe, however, that the slopes of the lines $x=\pm 4\sqrt{2}At$ approximate quite well the slopes of the boundaries of the wedge in the case of the AL lattice (panels (c) and (d)). It could be of interest to investigate if this fact could be another manifestation of the integrability of the AL lattice despite its discrete nature. In the outer sectors, the amplitude of $u_{n}(t)$ is $A_{*}$, of the undisturbed background, the only one which may sustain localized solutions in the case of the infinite lattice. On the other hand, since the numerical solutions are generated by solving numerically the periodic lattice, this background will never lose stability in terms of the amplitude in the outer regions since the solutions are initiated on the stable attractor of the finite system regarding its amplitude. Yet for the finite dimensional system, the MI dynamics in the wedge-shaped central region can be only transient as it is proved in Theorem III.1 and analyzed in the discussion of section V.1. Importantly, in the case of the finite lattice the wedge will eventually disappear as the system will select a stable wave number defining the global attractor, and the asymptotic state of the solution will be the plane wave of amplitude $A_{*}$ with the stable wavenumber. However, in subsections C and D that follow, we provide estimates for the $\ell^{2}_{per}$- distance between the solutions of the AL lattice and the DNLS (1.1). As we will discuss therein, these estimates provide additional theoretical justification for the proximity of the dynamics between the AL lattice and the DNLS (1.1), at least in the early stages of the MI behavior.